OpenStudy (anonymous):

(3x3 + 5x + 4) – (2x3 – 2x + 4)

OpenStudy (anonymous):

okay! We have to multiply the - to the other party , so we'll have : \[=3x^3 + 5x + 4 -2x^3+2x-4 \] Let's subtract : \[= x^3 + 7x\] let's simplify it and we'll get : \[= x(x^2+7)\] done ^_^

OpenStudy (anonymous):

thank you